# Afford2Act: Affordance-Guided Automatic Keypoint Selection for Generalizable and Lightweight Robotic Manipulation

### v1

@@ III-B Policy Design
At each interfere step $t$, taking the $K$ keypoints from keypoint distillation $\{p_{t,i}\}_{i=1}^{K}$, the question becomes: how do we architect a policy, $\pi_{\theta}$, that prioritizes the most task-relevant keypoints? After embedding each keypoint, $z^{1}=[\phi_{\theta}(p_{1})\dots\phi_{\theta}(p_{K})]^{\top}$, we then compute contextualized token features using a Transformer chunk, $z=\mathcal{T}_{\theta}(z_{1})$, similar to [31, 32] so that each keypoint is interpreted in the context of the others (e.g., a keypoint on the blade of a knife carries a different affordance than one on the handle). Since the number of keypoints $m$ can exceed the task-relevant cardinality, we introduce a gating network, $g_{\psi}$, that predicts a non-negative importance weight for each keypoint before pooling them into a single scene feature vector $h$:
$\hat{w}=g_{\psi}(z)=\operatorname{softmax}(w_{g}^{\top}z+b_{g}),\hat{z}=\hat{w}^{\top}z,\quad h=\frac{1}{K}\sum_{i=1}^{K}\hat{z}_{i}$
Then, the scene embedding, $h$ is passed through a shared action head that is later split into two linear layers: one for the delta actions and the other for the gripper state. While we use a two-layer MLP, the head is modular and can be replaced by a generative policy head [33, 34, 1].
@@ IV Experiments
    Keypoint Effectiveness: Can a policy trained exclusively on a compact set of 2D keypoints outperform policies trained on full images, dense features, and even a large vision-language foundation model?
@@
    Keypoint Priors: Which alternative keypoint selection strategies (e.g., using object bounding box, uniform grids of points, or manually selected points) can match the informativeness of our affordance-guided keypoints?
@@
    Generalization: How well does our policy generalize to new scenarios, including novel objects within seen categories, entirely new object categories, the presence of distractor objects, and other scene variations?
Additionally, we perform targeted ablation studies to test the robustness to different language prompts, to verify the necessity of the gating network, and to explore an extended two-stage pipeline for long-horizon tasks.
Figure 4: (Left) A figure of our physical setup, composing of a UR3e Robotic arm and one side RealSenseD435i Camera. (Right) shows the objects we used for seen (right) and unseen (left) scenarios.
Figure 5: Qualitative generalization on two tasks. Each row shows rollouts for Pour (top) and Kick (bottom). Left: training objects (seen). Right: held-out settings with unseen scenes (new backgrounds/tabletops) and, for Pour, unseen object instances. Within each panel we show three key frames (setup, approach, completion). Without any test-time adaptation, Afford2Act trained from single-view RGB using affordance-guided keypoints successfully transfers by attending to the functional part (mug handle / ball) while ignoring clutter.
@@ IV-A Setup and Tasks
Our method is driven by task semantics expressed as natural-language prompts that condition an affordance model. We focus on these core affordances—hold, cut_with, brush_with, pour, and kick—and systematically construct a variety of manipulation tasks by pairing each affordance with diverse objects and contexts. The mapping from affordances to example tasks is as follows:
Hold: Grasp and lift an object with a handle (e.g., a mug) using the hold affordance. (Fig. 1)
Cut: Perform a cutting motion on an object using a knife, guided by the cut_with affordance. (Fig. 3)
Brush: Sweep debris using a hand brush. This task leverages both the hold and brush_with affordances in coordination (holding the brush and using it to brush). (Fig. 2)
Pour: Execute a pouring motion with a vessel (e.g., tipping a cup or pitcher) using the pour affordance. (Fig. 5)
Stir: Stir the contents of a bowl using a spoon or similar tool, guided by a combination of hold (to grasp the spoon) and a stir affordance. (Fig. 1)
Kick: Push a ball into a goal using the kick affordance. 5
For each task, we train a policy and test it with a variety of seen and novel objects (see Fig. 4). The objects are physically diverse in size, shape, appearance, and material. This diversity allows us to evaluate generalization at both the instance level (new objects of a trained category) and the category level (objects from entirely new categories). We collected 40 teleoperated demonstrations per task for training, with object positions randomized across demonstrations.
IV-B Input Modality Baselines
We benchmark our approach against four baseline methods to evaluate the keypoints effectiveness:
  • RGB-D-BC: The framework as RGB-BC, but trained on 4-channel RGB-D images (RGB plus monocular depth from an Intel RealSense camera).
  • UAD-2D-BC: The same framework as RGB-BC, but the images are replaced with the affordance mask generated by UAD [3]. This baseline can verify the effectiveness of sparse keypoints rather than dense representations.
  • Foundation Model (Pi0): A large vision-language-action foundation model (Pi0) [35] fine-tuned on our demonstration data. We evaluate this model with a single side camera on a UR3e robot arm, which are settings that Pi0 has not encountered during its training.
Figure 6: Total Average success rate across input modalities. For each input modality, we test on 20 trials on each task, then average the results. We split the success rates of seen and unseen instances (detailed in Section IV-E).
IV-C Evaluation Protocol and Metrics
Success Rate: We evaluate each policy on 20 real-world episodes per task and report per-task success split by seen vs. unseen instances (Fig. 6). Our keypoint-based policy consistently outperforms RGB and RGB-D baselines, especially on unseen objects—likely because dense RGB inputs overfit to a single environment. For the UAD baseline, although it can extract dense features from the objects, it fail short with small number of demonstration because of its high-dimension representations. In contrast, the Pi0 foundation model struggled on most tasks. The first reason is that our setup uses only an external camera view, whereas Pi0 was designed to also use a wrist-mounted camera. Secondly, Pi0’s underlying vision-language model (PaLI-GEMMA) is relatively small, which required more data to refine and also made it sensitive to the language instructions.
Figure 7: Semantic Information rubric scores across input modalities on seen instances. For each input modality, we collected 40 trials for each task (we exclude the ball kick task as there is no notion of grasping). The success/failed shown as Y/N in the table for each sub-goal are manually recorded.
Semantic Trajectory Metrics: To diagnose failures beyond binary task success, we create a rubric to score each trajectory (as suggested by  [36]). The rubric breaks captures the success of several intermediate steps and not just the final task success (Fig. 7). In our evaluation, each sub-goal is checked independently, meaning we record whether the policy managed to achieve each individual step. To this end, we found cases where the RGB and RGB-D BC policies attempted to perform the final motion of the task (such as pouring or cutting) before actually grasping the object, indicating a confusion in the sequence of actions. In contrast, our keypoint-based policy typically reaches the final sub-goal, with occasional last-step slippage (object slipping from the gripper), suggesting strong early-stage competence.
Data Scaling: We study the effect of demonstration data size by training each method on progressively smaller subsets of the data. In particular, we downsampled the demonstrations to 50% and 25% of the full set and measured the success rates (Fig. 8. The results show that our keypoint-based method remains highly effective even with far fewer demonstrations, whereas the image-based baselines degrade significantly as the training data is reduced. This highlights the data efficiency of our approach compared to learning directly from raw images.
IV-D Keypoints Selections Baselines
@@
After demonstrating that a 2D keypoint-based policy can be highly effective, we ask whether our affordance-driven keypoint selection is essential. We compare our approach with several alternative keypoint selection strategies, each using $K=19$ (15 on object and 4 on end-effector) keypoints and the same downstream policy architecture for fairness:
@@
SAM [38] + DINO: Use the Segment Anything Model (SAM) to obtain segmentation masks, use YOLO to identify the target object’s mask, and then sample DINO correspondences within that segmented region.
@@
Hand-Picked (P3PO): An oracle baseline where an expert manually chooses the keypoints (following P3PO [6]), and these points are then tracked using DIFT + CoTracker.
@@
Fig. 9 summarizes the success rates of these strategies, averaged across all tasks, for both seen and unseen object instances. We find that the methods relying on YOLO detection do not significantly improve generalization to unseen object categories. This is likely because they depend on known object class labels (e.g., always looking for a “cup”), which can be misguiding when the object geometry changes (for example, encountering a teapot that might not be recognized as a cup). It also can suffer in cluttered scenes when the boxes are not tight or when multiple objects are present. In contrast, our affordance-driven approach does not depend on specific object labels; it directly identifies task-relevant regions (interaction hotspots) via the affordance model. We also note that the manually selected keypoints (P3PO) yield a success rate nearly on par with Afford2Act. However, our method achieves this performance without the tedious manual effort of choosing points, and it offers the flexibility of using natural language to specify the task.
IV-E Generalization Analysis
Beyond evaluating overall success, we systematically test four axes of generalization to assess how well Afford2Act transfers to new situations:
@@ IV-E Generalization Analysis
    Unseen Object (Same Category): Test the policy on novel instances of the same category (e.g., a new mug for the “hold” task or a different brush for “brush_with”).
@@
    Unseen Category: Apply the learned skill to an object from a category not encountered during training (e.g., using a spatula to stir instead of a spoon).
@@
    Background Variation: Evaluate the policy in environments with different backgrounds or lighting conditions.
@@
    Distractors: Introduce additional objects or even moving agents (e.g., a person performing an unrelated action nearby) to test the policy’s ability to stay focused on the target task in the presence of distractions.
Fig. 10 illustrates these generalization scenarios and the corresponding performance. Our keypoint-based policy exhibits strong transferability to novel object instances within seen categories and even to objects from entirely new categories (essentially achieving zero-shot generalization). It also remains robust under background changes and in cluttered environments. For example, in one challenging trial, the system successfully transferred the learned affordance to a teapot from a mug in a highly cluttered scene, correctly identifying the meaningful keypoints despite different geometry. Likewise, the policy proved resilient to dynamic distractions, even with a human actively moving nearby.
Figure 10: Top: Average success rate on four scenarios aggregated across all tasks. Bottom: Representative rollouts for each scenario. The small inset labeled “Ref” shows the training reference object for that task.
@@ IV-F Ablation Studies
We conducted additional ablations to highlight the contribution of different components of our pipeline:
@@
Language Prompt Robustness: We varied the wording of the affordance prompt to ensure our approach is not overly sensitive to specific language. Qualitatively, we found that synonyms or related phrases produce similar affordance keypoints. For example, using “sweep” instead of “brush_with” or “pick_up” instead of “hold” resulted in nearly identical affordance predictions on the relevant objects, indicating the model robustly understands the intended action.
@@
Gating Network: We removed the gating network so that each keypoint token receives equal weighting. This modification significantly degraded performance (especially in cluttered scenes), reducing the overall success rate from $86\%$ to $52\%$. During trajectory rollout, the gating mechanism allows the policy to focus only on the keypoints necessary.
@@
Effective Keypoint Count: We analyzed the gating network’s behavior by computing the effective number of keypoints being used by the policy at each time step. Given the gating output weights $w_{i}$} for the $K$ keypoint tokens, we define $K_{\text{eff}}=1/\sum_{i=1}^{K}w_{i}^{2}$. In our experiments, $K_{\text{eff}}$ was typically much smaller than the total number of points $K=19$, as shown in Fig. 11. This implies that the gating network effectively attends to only a subset of the keypoints at any given time, which helps the policy concentrate on the most informative features for the task.
@@
Figure 11: Average effective number of keypoints used (over all time steps) for each task. This data was collected through offline inference on all 40 training demos for each task.
IV-G Long-Horizon Skills via Two-Stage Re-Anchoring
For more complex, multi-phase tasks, we integrate Large Model task planning stage into our framework. We first use a vision-language model (VLM) planner to detect when the first sub-goal of the task has been completed (e.g., when the robot has successfully grasped the object). Once this condition is detected, we re-anchor half of the keypoints by providing an updated affordance prompt for the next phase of the task. In practice, this means replacing roughly half of the original keypoints with new ones tailored to the subsequent sub-task. This re-anchoring allows the policy to adapt to the changed state of the scene and effectively execute the next stage in the trajectory. We experimented with this augmentation on the Stir task, achieving an impressive $90\%$ success rate, further showing promise with incorporating VLMs with affordance-based keypoint selection (see video).

### v2

@@ III-B Policy Design
At each inference step $t$, we are given the $K$ distilled keypoints $\{p_{t,i}\}_{i=1}^{K}$. The policy must transform this unordered set into an action while prioritizing the keypoints that are most informative for the instructed manipulation. We first embed each keypoint into a token using a learnable embedding function $\psi_{\theta}$, forming a token set
$Z_{t}=[z_{t,1},\ldots,z_{t,K}]^{\top},\qquad z_{t,i}=\psi_{\theta}(p_{t,i}).$
We then compute contextualized token features with a Transformer encoder block: $\tilde{Z}_{t}=\mathcal{T}_{\theta}(Z_{t})$, so that each keypoint is interpreted in the context of the others (e.g., a point on a knife blade should be treated differently from a point on the handle) [11, 37].
Gating network. Keypoint distillation may return redundant or noisy candidates within the actionable region. We therefore introduce a gating network $g_{\theta}$ that assigns a non-negative importance weight to each token, allowing the policy to down-weight task-irrelevant keypoints. This explicitly separates candidate generation from relevance estimation within the control policy. Concretely, we compute per-keypoint scores and normalize them with a softmax:
$s_{t,i}=w_{g}^{\top}\tilde{z}_{t,i}+b_{g},\qquad\hat{w}_{t}=\operatorname{softmax}([s_{t,1},\ldots,s_{t,K}]),$
where $\hat{w}_{t}\in\mathbb{R}^{K}$ and $\hat{w}_{t,i}\geq 0$ with $\sum_{i=1}^{K}\hat{w}_{t,i}=1$. We form gated token features and aggregate them into a compact scene embedding: $h_{t}=\sum_{i=1}^{K}\hat{w}_{t,i}\,\tilde{z}_{t,i}$.
The resulting $h_{t}$ is then passed through a shared MLP action head, which branches into two linear outputs: one predicts the end-effector delta action and the other predicts the gripper command.
@@ IV Experiments
Figure 4: (Left) Our physical setup, consisting of a UR3e robotic arm and a side-mounted RealSense D435i camera. (Right) Objects used in seen (right) and unseen (left) scenarios.
Figure 6: Qualitative generalization on two tasks. Rows show rollouts for Pour (top) and Kick (bottom). Left: seen training objects. Right: held-out settings with unseen scenes (new backgrounds/tabletops) and, for Pour, unseen object instances. Each panel shows three key frames (setup, approach, completion). Without test-time adaptation, Afford2Act transfers by focusing on functional parts (mug handle/ball) while ignoring clutter.
@@
    Keypoint Priors (E1): How does our method compare against alternative keypoint selection strategies, and can our fully automatic pipeline match the performance of manually selected keypoints? (See Sec. IV-C.)
@@
    Keypoint Effectiveness (E2): Can a policy trained exclusively on a compact set of 2D keypoints outperform policies trained on full images or dense affordance features? (See Sec. IV-D.)
@@
    Generalization (E3): How well does our policy generalize to new scenarios, including novel objects within seen categories, new object categories (sharing the same functionality), the presence of distractor objects, and other scene variations? (See Sec. IV-E.)
Additionally, we perform targeted ablation studies to test robustness to different language prompts, verify the necessity of the gating network, and explore an extended two-stage pipeline for long-horizon tasks.
@@ IV-A Setup and Tasks
Our evaluation is organized around natural-language prompts that specify the intended interaction. We focus on a small set of affordances (hold, cut_with, sweep, pour, stir, kick) and instantiate them into six manipulation tasks: grasp-and-lift, cutting, brushing/sweeping (combining hold and sweep), pouring, stirring (combining hold and stir), and ball-kicking (Fig. 4). For each task, we train an independent policy and evaluate on both seen and novel objects. The object sets vary widely in size, geometry, appearance, and material, allowing us to measure transfer to new instances as well as to novel categories that share similar functional parts (Fig. 4).
IV-B Training and Evaluation Protocol
We train one policy per task using 40 teleoperated demonstrations and evaluate success on 20 real-world episodes per task, reported separately for seen and unseen objects. For keypoint-selection comparisons, all methods share the same downstream policy architecture and a fixed keypoint count of $K=19$ (15 object keypoints and 4 end-effector keypoints), so differences reflect representation quality rather than controller capacity. For semantic trajectory analysis, we annotate 40 seen-object trials per task, excluding ball-kicking because it has no grasp phase, and for data-scaling experiments, each method is retrained on reduced demonstration subsets. This efficiency comes from running pretrained affordance and correspondence modules only on the first frame, then training the policy on sparse 2D keypoint trajectories. In our implementation, training one Afford2Act policy from 40 demonstrations takes approximately 15 minutes on a single RTX 2080 GPU.
IV-C Keypoint Priors (E1): Baseline Comparisons
@@
To demonstrate the strength of affordance-guided keypoint selection, we compare our approach with several alternative keypoint selection strategies, each using $K=19$ (15 on the object and 4 on the end-effector) keypoints and the same downstream policy architecture for fairness:
@@
SAM [15] + DINO: Use the Segment Anything Model (SAM) to obtain segmentation masks, then use YOLO to identify the target object’s mask and sample DINO correspondences within that segmented region [29].
@@
Hand-Picked (P3PO-style): A manual expert-selected baseline where keypoints are specified per task (following P3PO [17]) and then tracked using DIFT + CoTracker.
@@
Fig. 5 summarizes the success rates of these strategies, averaged across all tasks, for both seen and unseen object instances. We find that methods relying on YOLO detection do not significantly improve generalization to unseen object categories. This is likely because they depend on known object class labels (e.g., always looking for a “cup”), which can be misleading when object geometry changes (e.g., encountering a teapot that might not be recognized as a cup). They can also suffer in cluttered scenes when boxes are not tight or when multiple objects are present. In contrast, our affordance-driven approach does not depend on specific object labels; it directly identifies task-relevant regions (interaction hotspots) via the affordance model. The manual expert-selected baseline serves as an upper-bound reference for semantically meaningful sparse keypoints. Afford2Act reaches comparable performance without per-task point specification, replacing manual geometric priors with a fully automatic, language-conditioned procedure.
IV-D Keypoint Effectiveness (E2): Input Modality Baselines
To showcase the strength of keypoints over dense inputs, we conduct an extensive benchmark of our approach against three baselines across success rate, semantic trajectory metrics, and data scaling:
  • UAD-2D-BC: The same framework as RGB-BC, but the input images are replaced with affordance masks generated by UAD [26].
  • RGB-D-BC: The same framework as RGB-BC, but trained on 4-channel RGB-D images (RGB plus monocular depth from an Intel RealSense camera).
We also evaluated the Pi0 foundation model [3], but did not obtain any successful rollouts in our setting. This result is expected given the mismatch in assumptions: our experiments use a single external camera, whereas Pi0 is designed to leverage additional wrist-mounted viewpoints. In addition, our robot’s embodiment (UR3e) is not represented in Pi0’s training distribution [28].
Figure 7: Total average success rate across input modalities. For each input modality, we run 20 trials per task and then average the results. We report seen and unseen performance separately (detailed in Section IV-E).
Evaluation Metrics. Success Rate: We evaluate each policy on 20 real-world episodes per task and report per-task success rates split by seen vs. unseen instances (Fig. 7). Our keypoint-based policy consistently outperforms RGB and RGB-D baselines, especially on unseen objects, because dense pixel inputs tend to overfit to a single environment. For the UAD baseline, although it can extract dense features from objects, it falls short with a small number of demonstrations because of its high-dimensional representations.
Figure 8: Semantic information rubric scores across input modalities on seen instances. For each input modality, we collected 40 trials per task (excluding ball-kicking, which has no grasping sub-goal). Success/failure for each sub-goal is shown as Y/N and was recorded manually.
Semantic Trajectory Metrics: To diagnose failures beyond binary task success, we create a rubric to score each trajectory (as suggested by [16]). The rubric captures success on several intermediate steps, not just final task completion (Fig. 8). In our evaluation, each sub-goal is checked independently, meaning we record whether the policy achieves each individual step. We found cases where the RGB and RGB-D BC policies attempted the final motion of the task (such as pouring or cutting) before grasping the object, indicating confusion in action sequencing. In contrast, our keypoint-based policy typically reaches the final sub-goal, with occasional last-step slippage (object slipping from the gripper), suggesting strong early-stage competence.
Data Scaling: We study the effect of demonstration data size by training each method on progressively smaller subsets of the data. In particular, we downsampled the demonstrations to 50% and 25% of the full set and measured the success rates (Fig. 9). The results show that our keypoint-based method remains highly effective even with far fewer demonstrations, whereas the image-based baselines degrade significantly as the training data is reduced. This highlights the data efficiency of our approach compared to learning directly from raw images.
IV-E Generalization (E3): Analysis
Beyond overall success rates, we evaluate transfer along four complementary axes to characterize how Afford2Act behaves under distribution shift:
@@ IV-E Generalization (E3): Analysis
    Unseen Instance (Same Category): Evaluate on novel instances from the training category (e.g., a different mug for hold or a different brush for brush_with).
@@
    Novel Category (Shared Functional Parts): Evaluate on objects from categories not seen during training that support the same interaction (e.g., stirring with a spatula instead of a spoon).
@@
    Scene Variation: Evaluate under changes in background and lighting.
@@
    Distractors: Introduce additional objects and moving agents (e.g., a person performing an unrelated action nearby) to test robustness to task-irrelevant dynamics.
Fig. 10 illustrates these generalization scenarios and the corresponding performance. Our keypoint-based policy exhibits strong transfer to novel object instances within seen categories and also transfers to several new categories in our test scenarios, particularly when objects share similar functional parts, suggesting promising zero-shot generalization. It also remains robust under background changes and in cluttered environments. For example, in a particularly challenging trial, a policy trained to lift a mug successfully lifted a teapot in a heavily cluttered scene, with the pipeline still localizing the correct interaction keypoints despite substantial changes in object geometry and background. Likewise, the policy proved resilient to dynamic distractions, even with a human actively moving nearby.
Figure 10: Representative rollouts for each scenario. The small inset labeled “Ref” shows the training reference object for that task. We report the success rate for each scenario across all experiments in parentheses.
@@ IV-F Ablation Studies
We conducted additional ablations to highlight the contributions of different components of our pipeline:
@@
Language Prompt Robustness: We varied the wording of the affordance prompt to ensure our approach is not overly sensitive to specific language. Qualitatively, we found that synonyms or related phrases produce similar affordance keypoints. For example, using “brush_with” instead of “sweep” or “mix” instead of “stir” resulted in nearly identical affordance predictions on the relevant objects, indicating that the model robustly captures the intended action semantics.
@@
Gating Network: We removed the gating network so that each keypoint token receives equal weight. This modification significantly degraded performance (especially in cluttered scenes), reducing the overall success rate from $86\%$ to $52\%$. During trajectory rollout, the gating mechanism allows the policy to focus only on the necessary keypoints.
@@
Effective Keypoint Count: We analyzed the gating network’s behavior by computing the effective number of keypoints used by the policy at each time step. Given the gating output weights $w_{i}$} for the $K$ keypoint tokens, we define $K_{\text{eff}}=1/\sum_{i=1}^{K}w_{i}^{2}$. In our experiments, $K_{\text{eff}}$ was typically much smaller than the total number of points $K=19$, as shown in Fig. 11. This implies that the gating network effectively attends to only a subset of the keypoints at any given time, which helps the policy concentrate on the most informative features for the task.
@@
Figure 11: (a) Average effective number of keypoints used (over all time steps) for each task. This data was collected through offline inference on all 40 training demos for each task. (b) Failure case due to weak affordance masking, leading to the gripper colliding with the spatula.
IV-G Two-Stage Re-Anchoring for Multi-Phase Tasks
To examine whether affordance-grounded keypoints can extend beyond single-stage interactions, we conduct a preliminary study on the Stir task using two-stage re-anchoring. This task is explicitly multi-phase: the robot must first grasp the spoon, then move to the bowl, and finally execute the stirring motion. A vision-language model (VLM) detects completion of the initial grasping phase; after this trigger, we update the affordance prompt and re-anchor a subset of keypoints for the second phase. In this setting, the approach achieves a $90\%$ success rate, indicating that online keypoint refresh can be useful for multi-phase manipulation. Future work will evaluate this mechanism across additional multi-phase tasks, alternative phase-transition criteria, and longer-horizon interactions.
IV-H Failure Cases and Limitations
A primary failure mode arises when the affordance prior is inaccurate or incomplete. Because keypoint distillation is conditioned on the predicted actionable region, an imperfect mask can omit the true contact area or include distractors, leading to unstable keypoints and incorrect control (Fig. 11). A second limitation is partial observability from a single RGB view: heavy occlusions or viewpoint ambiguity can break affordance localization and tracking even when the behavior is feasible. These failures mainly reflect perception reliability rather than the keypoint-based policy design, and could be mitigated with confidence-aware filtering, lightweight re-anchoring, and occlusion/mask-noise augmentation.